# VisPhyWorld: Probing Physical Reasoning via Code-Driven Video Reconstruction

### v1

@@ 4 Experiments
Second, the choice of code backend affects reconstruction quality. Across LLMs, Three.js variants yield better perceptual reconstruction than their P5.js counterparts, as reflected by lower LPIPS in most pairs, despite sharing the same conditioning inputs and prompt. Concretely, for GPT-5, switching to Three.js reduces LPIPS error by nearly 40% ($0.29\rightarrow 0.17$) and boosts structural similarity (SSIM) from $0.74$ to $0.94$. Visually, this corresponds to better preservation of object identity, as illustrated in Figure 1. Since the physics engine is fixed, this performance gap confirms that the simulator’s expressivity affects the model’s ability to ground visual evidence. In other words, program structure and simulator interface materially affect how well a model can translate visual evidence into a stable physical hypothesis.
Third, pixel-space baselines exhibit a complementary profile: they can score competitively on some frame-feature semantics, but their failures are harder to attribute to specific physical causes, such as friction, restitution, or contact timing, since the generation process does not expose interpretable latent variables. Veo-3.1 attains reasonable semantic similarity, for example reaching DINO $\sim 0.88$, yet it does not expose an explicit simulator state for diagnosis or controlled interventions and often exhibits deficiencies in physical understanding by producing trajectories with implausible motion or contact events(see Sec. 4.2). Conversely, our code-driven approach maintains competitive semantic and motion scores while exposing executable states; e.g., GPT-5 (threejs) achieves DINO $0.8556$ with RAFT-EPE $33.6473$. This enables controlled interventions (e.g., varying friction/mass while holding layout fixed) that can isolate whether an error originates from object discovery, state initialization, or contact modeling, aligning with our goal of turning “physics understanding” into a testable, executable object.
Finally, we report holistic quality using a Gemini-2.5-Pro judge (see Appendix C.6), which aggregates multiple perceptual and physical cues into a single preference signal. This holistic score aligns with strong visual alignment for some backends. For example, Gemini-3-Pro (threejs) reaches the highest Gemini score, $3.80$, while penalizing visibly implausible or degraded generations. For instance, Qwen3-VL-Plus (p5js) scores $1.46$ alongside poor perceptual/semantic alignment. This judge complements the automatic metrics by capturing visually salient failure modes (e.g., missing motion, implausible contacts) that may not be fully reflected in any single metric family. Together, these results indicate that VisPhyBench and our metric suite jointly provide a multi-view, diagnostic measurement of LLM visual and physical competence under executable simulation.
@@ 4.2 Case Study
While VisPhyWorld shows promising results on physics-aware video generation and evaluation, it has several limitations. First, our experiments are conducted on synthetic, simulator-driven scenes with controlled object layouts and camera motion, so generalization to high-resolution, in-the-wild videos remains untested. Fundamentally limited by the capabilities of current MLLM and the complexity of modern engines, VisPhyWorld can reliably generate code only for relatively simple rigid-body scenes: although we experimented with large engines such as Unreal Engine and Blender, we found that, without human intervention, existing MLLMs cannot, within a small fixed number of calls, autonomously produce and repair simulation code to render a stable, visually plausible video in these more complex environments. Finally, we currently target relatively short clips with moderate motion complexity, and do not explicitly address long-horizon interactions, complex 3D reasoning, or stylized or heavily cluttered scenes, which we leave for future work. Future work could integrate stronger 3D perception for scene initialization and agentic workflows with domain-specific fine-tuning.

### v2

@@ 4 Experiments
Second, the choice of code backend affects reconstruction quality. Across LLMs, Three.js variants yield better reconstruction than their P5.js counterparts, as reflected by lower LPIPS in most pairs, despite sharing the same conditioning inputs and prompt. Concretely, for GPT-5, switching to Three.js reduces LPIPS error by nearly 40% ($0.29\rightarrow 0.17$) and boosts structural similarity (SSIM) from $0.74$ to $0.94$. Visually, this corresponds to better preservation of object identity, as illustrated in Figure 1. Since the physics engine is fixed, this performance gap confirms that the simulator’s expressivity affects the model’s ability to ground visual evidence. In other words, program structure and simulator interface materially affect how well a model can translate visual evidence into a stable physical hypothesis.
Third, pixel-space baselines exhibit a complementary profile: they can score competitively on some feature semantics, but their failures are harder to attribute to specific physical causes, such as friction, restitution, or contact timing, since the generation process does not expose interpretable latent variables. Veo-3.1 attains reasonable semantic similarity, for example reaching DINO $\sim 0.88$, yet it does not expose an explicit simulator state for diagnosis or controlled interventions and often exhibits deficiencies in physical understanding by producing trajectories with implausible motion or contact events(see Sec. 4.2). Conversely, our code-driven approach maintains competitive semantic and motion scores while exposing executable states; e.g., GPT-5 (threejs) achieves DINO $0.8556$ with RAFT-EPE $33.6473$. This enables controlled interventions (e.g., varying friction/mass while holding layout fixed) that can isolate whether an error originates from object discovery, state initialization, or contact modeling, aligning with our goal of turning “physics understanding” into a testable, executable object.
Finally, we report holistic quality using a Gemini-2.5-Pro judge (see Appendix C.6), which aggregates multiple perceptual and physical cues into a single preference signal. This holistic score aligns with strong visual alignment for some backends. For example, Gemini-3-Pro (threejs) reaches the highest Gemini score, $3.80$, while penalizing visibly implausible or degraded generations. For instance, Qwen3-VL-Plus (p5js) scores $1.46$ alongside poor perceptual/semantic alignment. This judge complements the automatic metrics by capturing visually salient failure modes (e.g., missing motion, implausible contacts) that may not be fully reflected in any single metric family. Together, these results indicate that VisPhyBench and our metric suite provide a multi-view, diagnostic measurement of LLM visual and physical competence under executable simulation.
@@ 4.2 Case Study
While VisPhyWorld shows promising results on physics-aware video generation and evaluation, it has limitations. First, our experiments are conducted on synthetic, simulator-driven scenes with controlled object layouts and camera motion, so generalization to high-resolution, in-the-wild videos remains untested. Fundamentally limited by the capabilities of current MLLMs and the complexity of modern engines, VisPhyWorld can reliably generate code only for relatively simple rigid-body scenes: although we experimented with large engines such as Unreal Engine and Blender, we found that, without human intervention, existing MLLMs cannot, within a small fixed number of calls, autonomously produce and repair simulation code to render a stable, visually plausible video in these more complex environments. Finally, we currently target relatively short clips with moderate motion complexity, and do not address long-horizon interactions, complex 3D reasoning, or stylized or heavily cluttered scenes, which we leave for future work. Future work could integrate stronger 3D perception for scene initialization and agentic workflows with domain-specific fine-tuning.